datal (0.9 1.32 1.96 1.85 2.29 1.42 1.35 147 1.74 1.82 1.3 147 192) data2 (1.65 2.06 1.55 1.95 1.35 1.78 2.14 1.63 1.66 1.05 1.71 1.27)

Given the following 2 sets of experimental data find the sample mean, sample standard deviation, and sample variance using formulas shown in class. Then, find a histogram for the range of 0.6 to 2.4 with intervals of width 0.2. Finally, plot a normal distribution with the calculated mean and standard deviation. Hint: You must start by putting the two data sets into a single vector.
Answer #1

R code:
To get til histogram

data1=c(0.9,1.32,1.96,1.85,2.29,1.42,1.35,1.47,1.74,1.82,1.3,1.47,1.92)
data2 = c(1.65,2.06,1.55,1.95,1.35,1.78,2.14,1.63,1.66,1.05,1.71,1.27)
data3<- c(data1,data2)
data3
mean(data3)
sd(data3)
class(data3)
hist(data3,breaks=c(0.9,seq(0.6,2.4,0.2)))

output:

data3
[1] 0.90 1.32 1.96 1.85 2.29 1.42 1.35 1.47 1.74 1.82 1.30 1.47 1.92 1.65 2.06 1.55 1.95 1.35 1.78
[20] 2.14 1.63 1.66 1.05 1.71 1.27
> mean(data3)
[1] 1.6244
> sd(data3)
[1] 0.3393877
> class(data3)
[1] "numeric"
> hist(data3,breaks=c(0.9,seq(0.6,2.4,0.2)))

mean=1.6244

standard deviation=0.3393877

Histogram of data3 1.5 2.0 1.0 data3